OpenStudy (anonymous):

A piano tuner strikes a tuning fork for note A above middle C and sets in motion vibrations that can be modeled by the equation y = 0.001 sin 880 Pi(T). Find the amplitude and period for the function.

OpenStudy (anonymous):

period of \(\sin(bx)\) is \(\frac{2\pi}{b}\) so your period is \[\frac{2\pi}{800\pi}=\frac{1}{400}\]

OpenStudy (anonymous):

amplitude is the number outside of the sine, i.e. \(0.001\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw